Two boxes, each with mass 20 kg, are stacked one on top of the other. Alex lifts the boxes at a constant speed, pushing upward on the bottom of the lower box.

a) How much force must Alex apply to lift the two boxes?

b) How much force does the bottom box exert on the top box?

c) How much force does the top box exert on the bottom box?

d) How much force does the bottom box exert on Alex?
Answer #1

here,

mass of each box , m = 20 kg

a)

the force must Alex apply to lift the two boxes , Fa = 2 * m * g

Fa = 2 * 20 * 9.81 N = 392.4 N

b)

the force does the bottom box exert on the top box , Fb = m * g = 20 * 9.81 = 196.2 N

c)

the force does the top box exert on the bottom box, Fc = 196.2 N downwards

d)

the force does the bottom box exert on Alex , F = 2 mg downwards

F = 392.4 N downwards